OpenStudy (anonymous):

Give the measure of one interior angle of a regular decagon A.120 B. 135 C. 144 D. 100

OpenStudy (mathstudent55):

Do you know of a formula to find the sum of the measures of the interior angles of a polygon?

OpenStudy (anonymous):

((sides-2)*180)/sides = angle measure

OpenStudy (anonymous):

what are you looking for ?

OpenStudy (anonymous):

The sum of the interior angles can be found by using this formula : 180(n - 2)....n being the number of sides. a decagon has 10 sides... 180(10 - 2) = 180(8) = 1440 (this is the total of the angles) Each interior angle can be found with this formula : (180(n - 2)) / n (180(10 - 2)) / 10 (180(8)) / 10 1440/10 = 144